• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 23:53:12 on January14,2003

    Entry number 91922

    keyword=Swing shift summary

    4:20 left arm Q2 requires access.
    5:05 beam back
    6:02 Beam off to improve tune
    6:12 beam back
    6:30 Beam back after more tuning
    7:05 Beam down to cycle beamline magnets
    7:20 Compton group decides to delay measurement till tomorrow
    8:20 NMR/EPR measurement begins
    8:32 beam back on target
    10:25 Completed kinematic 4.7.3.
    --Began field rotation to 180 deg.
    ----Target half wave plate in
    10:35 Began kinematic 4.7.5